# Understanding Ethereum Mempool Security under Asymmetric DoS by Symbolized Stateful Fuzzing

### v3

@@ 5.1.1 State Search Algorithm
6:             emit($ops^{\prime}$, $st^{\prime}$, "Found an exploit");
@@
The core data structure is a seed corpus or $sdb$ which stores a list of input-state pairs or seeds. Upon running the algorithm, the corpus maintains the mempool states covered so far by the algorithm execution. The algorithm runs an outer loop that continues until the corpus is empty, or only the states of zero energy are left, or timeout. In each iteration, the algorithm retrieves the next seed from the corpus based on how promising the seed is in reaching a state triggering the oracle (a.k.a., the energy [22]). A seed consists of a symbolized input $ops$, which consists of transaction symbols, and a symbolized state $st$ reached by running a transaction sequence instantiated from the input $ops$ (Line 3). The algorithm runs an inner loop, in each iteration of which it mutates the input in the current seed ($\langle{}st,in\rangle{}$) and executes the mutated input against a reinitialized empty mempool, producing end state $st^{\prime}$ (Line 4). If the end state satisfies ADAMS conditions (i.e., the test oracle), it emits the mutated input-state pair $ops^{\prime},st^{\prime}$ as a newly found exploit. The algorithm further checks the feedback: The feedback is positive if the mutated input $ops^{\prime}$ brings the mempool state $st^{\prime}$ closer to triggering the test oracle than $st$. This entails mutated state $st^{\prime}$ to be different from state $st$ (i.e., increased state coverage) and state $st^{\prime}$ to achieve larger damage or lower attack cost. In case of positive feedback, the algorithm would add mutated input-state pair $ops^{\prime},st^{\prime}$ to the corpus.
@@
Estimate state cost: When mpfuzz determines how promising a state is, it needs to look beyond the current state and into all possible descendant states. We propose a heuristic that optimistically estimates the descendant-state costs given a current state. The key intuition is the following: In Ethereum, the validity of a child transaction (under symbol $\mathcal{C}$) depends on its parent/ancestor transaction. Thus, even though a transaction in the current state is valid, the transaction can be “turned” into an invalid one in subsequent states. We thus attribute, optimistically, the cost of transaction $\mathcal{C}$ to value $1$, so that it is preferable to a parent transaction $\mathcal{P}$ in input mutation and seed selection, and turning $\mathcal{C}$ into an invalid one like $\mathcal{L}$ or $\mathcal{F}$ is also encouraged. The cost profiles used in mpfuzz are summarized in Table 2.
@@ 6 Found Exploits
Table 3: ADAMS exploit patterns found by mpfuzz across Ethereum clients; $XT_{1-7}$ are eviction based, and $XT_{8-9}$ are locking based. $XT_{1-3}$ are known patterns in DETER [28], while others are new. ✗ indicates the presence of a bug, indicates the fixing of a bug after our reporting, and ✓ indicates the fixing of a bug by previous works.
@@
On the six public-transaction clients, mpfuzz leads to the discovery of $22$ bugs, as listed in Table 3, including $7$ conforming to the known DETER attacks on clients of historical versions and $15$ new bugs on the clients of the latest versions.
Other clients, including the three PBS clients and three clients on Ethereum-like networks, are mostly forks of the Geth clients except for minor code changes [^4] For instance, go-opera adopts its own fix against $XT_{1}$ based on the fork of Geth $<v1.11.4$., and on them, mpfuzz discovered $13$ bugs of a similar nature to those found on Geth (of the historical and latest versions), as seen in Table 3.
@@
6.1 Found Eviction Attacks
Exploit $XT_{1}$: Direct eviction by future transactions: In this attack, given a mempool’s initial state storing normal transactions, the attacker sends future transactions at high GasPrice to evict the normal transactions. This exploit is essentially the DETER-X attack [28].
In practice, Geth ($\leq{}v1.11.4$), Nethermind, Besu, and EigenPhi are vulnerable under this exploit, as in Table 3.
Exploit $XT_{2}$: Direct eviction by latent overdraft transactions: In this eviction attack, the attacker sends latent-overdraft transactions at high GasPrice to evict the normal transactions initially stored in the target mempool. These transactions are sent from $k$ accounts, each of which sends $l$ transactions, denoted by $k\times{}l$. The intention is to evade the limit on the number of transactions per sender. The evasion increases the attacker cost from one pending transaction to multiple. This exploit is essentially the DETER-Z attack [28].
@@ 6.1 Found Eviction Attacks
Exploit $XT_{4}$: Indirect eviction by valid-turned-overdraft transactions: This eviction attack works in two steps: 1) The attacker first sends valid transactions at high GasPrice to evict normal transactions initially stored in the mempool. These transactions are sent from $k$ accounts, each of which sends $l$ transactions. 2) She then sends $k$ transactions; each of the transactions is of nonce $1$, at a high GasPrice, of high value $v_{1}$, and from the same $k$ accounts in Step 1). The transactions would replace the transaction of the same sender and nonce sent in Step 1). Once the replacement is finished, they turn their child transactions into latent overdraft. Specifically, if a sender’s balance is $bal$ and the Ether value of transaction of nonce 2 is $v_{2}$, then $v_{1}$ is carefully crafted to enable turned latent overdraft, that is, $v_{1}<bal$ and $v_{1}+v_{2}>bal$.
@@
Exploit $XT_{6}$ works in three steps: 1) it first evicts the initial mempool of normal transactions with $m^{\prime}/py_{2}^{\prime}=384$ sequences, each of $16$ transactions sent from one unique sender, 2) it then sends $py_{1}^{\prime}/(py_{2}^{\prime}-1)=65$ transactions to evict the parent transactions and creates $py_{1}^{\prime}=1024$ future transactions in the mempool, 3) it sends one sequence of $5120$ transactions from one sender to evict the normal transactions sent in Step 1), and 4) it sends one transaction to evict the parent transaction in Step 3, leaving the mempool of just one valid transaction. Step 3) can succeed because the precondition (w.r.t. $py_{3}^{\prime}=5120$) of limiting transactions of the same sender does not hold.
@@
The mempool on Nethermind $<v1.18.0$ can be reversible: Minimally, suppose a two-slot mempool stores $tx_{1}$ of low GasPrice from one sender and $tx_{2}$ of medium GasPrice from another sender $A2$ and receives $tx_{3}$ as a child of $tx_{1}$ and of high GasPrice. For instance, $tx_{1}\begin{bmatrix}{A}&{1}\\ {1}&{1}\end{bmatrix}$, $tx_{2}\begin{bmatrix}{B}&{1}\\ {1}&{3}\end{bmatrix}$, $tx_{3}\begin{bmatrix}{A}&{1}\\ {2}&{5}\end{bmatrix}$. The Nethermind $<v1.18.0$ mempool would admit $tx_{3}$ and evict $tx_{2}$. After that, if one resends the evicted $tx_{2}$ back to the mempool storing $tx_{1}$ and $tx_{3}$, the Nethermind $<v1.18.0$ mempool would admit $tx_{2}$ and evict $tx_{3}$, looping back to its initial state.
An attacker observing reversible mempool mounts Exploit $XT_{7}$ to evict the mempool while bringing down attack costs. On Nethermind $<v1.18.0$, the attack works in two steps: 1) She first sends $k$ transaction sequences from $k$ senders, [^5] Nethermind does not limit transactions per sender, thus $k=1$ in actual attacks. each of $l$ transactions (i.e., $k\times{}l$). In each $l$-transaction sequence, the nonce-$1$ transaction has a low GasPrice, say $f_{1}$, while all its child transactions are of high GasPrice, evicting normal transactions initially stored in the mempool. 2) The attacker sends $k\times(l-1)$ transactions from $k\times(l-1)$ new senders, all with GasPrice slightly higher than $f_{1}$. The mempool would admit these transactions to the mempool by evicting all child transactions sent in Step 1), bringing down the attack costs.
6.2 Found Locking Attacks
@@
Exploit $XT_{9}$: Locking mempool of no turning: Because of the risk of evicting a parent transaction that could lead to turning, certain mempool is designed to restrict eviction victims to child transactions; that is, only the transaction of maximal nonce under a given sender can be evicted. For instance, given an arriving transaction $tx$, the OpenEthereum mempool finds an existing child transaction $tx^{\prime}$ with a lower GasPrice than $tx$ and evicts it to admit $tx$.
In Exploit $XT_{9}$, the attacker observes empty slots in the mempool (e.g., created by arriving blocks) and sends the same number of transactions to occupy them. The transactions are sent from one account where the child transaction has a higher GasPrice $f_{1}$ than normal transactions, while all its parent transactions have minimal fees. A normal transaction that arrives subsequently is declined because $f_{1}$ is higher than a normal transaction. Because each of its parent transactions has a low GasPrice, the mempool is locked at a low cost.
6.3 Evaluation of Found Turning Exploits
This and next subsections show the evaluation of turning attacks and locking attacks, and due to the space limit, the evaluation of other ADAMS attacks is deferred to Appendix D.
@@ 6.3.1 Evaluation in Testnet
To monitor transaction propagation, we launch an independent supernode (called measurement node) in the same testnet. The measurement node runs a reconfigured Geth client where the limit of peers/neighbors is removed and can connect to as many neighbors as possible. The measurement node is not (directly) connected to the attacker node. When setting up our experiments, we ran the measurement node in Goerli for seven days and found the node is stabilized at $290$ neighbors.
@@
6.4 Evaluation of Found Locking Exploits
6.4.1 Exploit Extension
@@
Actual exploit $XT_{8a}$: In $XT_{8}$, the attack cost increases with the block $BasePrice$. We manually propose a method to decrease the block $BasePrice$ in Ethereum by mounting an eviction attack on the previous block. Specifically, in Ethereum (after EIP1559), given a recently produced block $i$, the block $BasePrice$ of block $i+1$ is calculated dynamically as follows:
@@
Therefore, if an eviction attack can persistently lower the Gas utilization of recent blocks, the current $BasePrice$ can be reduced, which decreases the costs of locking attacks on the current block. In Exploit $XT_{8a}$, we mount a series of eviction attacks first and then a locking attack against the Reth node.
Specifically, suppose in a network of a Reth node and a block validator node, the $XT_{8a}$ attacker first keeps sending eviction attacks directly to the validator node for several consecutive blocks. When observing the block $BasePrice$ drop sufficiently, the attacker mounts the regular locking attack $XT_{8}$ against the Reth node.
6.4.2 Attack Evaluation on Reth
@@
Figure 5: Evaluation of locking attack $XT_{8}$ on Reth
@@
Evaluation of $XT_{8}$: To evaluate $XT_{8}$, we set up the workload generator that sends normal transactions to the validator Geth node. The non-validator Reth node has not yet joined the network. We let the attacker node first send eviction attacks ($XT_{6}$) directly to the victim validator for $35$ consecutive blocks until it observes the block $BasePrice$ drops sufficiently to $1$ Gwei, which occurs at the $45$-th in our experiment. The Reth node then joins the network with an empty mempool. The attacker node mounts a locking attack $XT_{8}$ with transactions of price $5$ Gwei to occupy the Reth node’s mempool.

### v4

@@ 5-A1 State Search Algorithm
6:             emit($ops^{\prime}$, $st^{\prime}$, ”Found an exploit”);
@@
The core data structure is a seed corpus or $sdb$ which stores a list of input-state pairs or seeds. Upon running the algorithm, the corpus maintains the mempool states covered so far by the algorithm execution. The algorithm runs an outer loop that continues until the corpus is empty, or only the states of zero energy are left, or timeout. In each iteration, the algorithm retrieves the next seed from the corpus based on how promising the seed is in reaching a state triggering the oracle (a.k.a., the energy [36]). A seed consists of a symbolized input $ops$, which consists of transaction symbols, and a symbolized state $st$ reached by running a transaction sequence instantiated from the input $ops$ (Line 3). The algorithm runs an inner loop, in each iteration of which it mutates the input in the current seed ($\langle{}st,in\rangle{}$) and executes the mutated input against a reinitialized empty mempool, producing end state $st^{\prime}$ (Line 4). If the end state satisfies ADAMS conditions, it emits the mutated input-state pair $ops^{\prime},st^{\prime}$ as a newly found exploit. The algorithm further checks the feedback: The feedback is positive if the mutated input $ops^{\prime}$ brings the mempool state $st^{\prime}$ closer to triggering the test oracle than $st$. This entails mutated state $st^{\prime}$ to be different from state $st$ (i.e., increased state coverage) and state $st^{\prime}$ to achieve larger damage or lower attack cost. In case of positive feedback, the algorithm would add mutated input-state pair $ops^{\prime},st^{\prime}$ to the corpus.
@@
Estimate state cost: When mpfuzz determines how promising a state is, it needs to look beyond the current state and into all possible descendant states. We propose a heuristic that optimistically estimates the descendant-state costs given a current state. The key intuition is the following: In Ethereum, the validity of a child transaction (under symbol $\mathcal{C}$) depends on its parent/ancestor transaction. Thus, even though a transaction in the current state is valid, the transaction can be “turned” into an invalid one in subsequent states. We thus attribute, optimistically, the cost of transaction $\mathcal{C}$ to value $1$, so that it is preferable to a parent transaction $\mathcal{P}$ in input mutation and seed selection, and turning $\mathcal{C}$ into an invalid one like $\mathcal{L}$ or $\mathcal{F}$ is also encouraged. The cost profiles used in mpfuzz are summarized in Table II.
5-B Exploit Extension
The initial short exploit identified on MUT necessitates further automatic development to create a more extensive exploit that is effective within a larger mempool. The key idea in extending the small exploit to a more comprehensive one is to ensure the same admission event observed in the smaller MUT. We first define the admission event as given an arriving transaction $tx_{i}$ at a mempool of initial state $st_{0}$, the admission of $tx_{i}$ transitions the mempool into an end state $tx_{n}$. The admission event is denoted by $f(st_{0},tx_{i})\Rightarrow{}st_{n}$. We extract the admission event patterns, representing a state transition from the initial state to the end state. In the pattern, we only symbolize the mempool slots that are updated during the state transition. A pattern is denoted as $\langle{}st_{0}^{\prime},tx_{i},st_{n}^{\prime}\rangle{}$. For example, in the state transition from $\mathcal{NNPC}$ to $\mathcal{NPPC}$ by admitting a transaction $\mathcal{P}_{0}$, the symbolized state of a mempool slots transitions from $\mathcal{N}$ to $\mathcal{P}$. Thus, we extract the pattern as $\langle{}\mathcal{N},\mathcal{P},\mathcal{P}\rangle{}$.
we extract all the unique admission event patterns in the small exploit and store them into an admission event pattern set ($aps$). We then construct the attack transaction sequence on the larger-size mempool by aligning the patterns identified from the small exploit with corresponding admission events. This entails validating that the admission event pattern of a given event $f(st_{0},tx_{i})\Rightarrow{}st_{n}$ on the larger-size mempool exists within the $aps$. Specifically, mpfuzz first selects $ap_{i}$ from the set $aps$ whose initial state $st_{0}$ matches the state of the large mempool. We then instantiate $tx_{i}$ of the $ap_{i}$ and send it to the large mempool. If the state transition of the large mempool $\langle{}st_{0}^{\prime},tx_{i},st_{n}^{\prime}\rangle{}$ aligns with the pattern $ap_{i}$, this positive outcome prompts mpfuzz to proceed with the extension process accordingly. Otherwise, mpfuzz backtracks by rolling back the mempool state to its state prior to the admission of $tx_{i}$ and tries the next pattern. mpfuzz tries the patterns in the decreasing order of $opcost()$, which means if multiple admission event patterns in $aps$ yield positive results, mpfuzz proceeds the extension process by applying the pattern with the lower $opcost()$.
@@ 6 Found Exploits
TABLE III: ADAMS exploit patterns found by mpfuzz across Ethereum clients; $XT_{1-7}$ are eviction based, and $XT_{8-9}$ are locking based. $XT_{1-3}$ are known patterns in DETER [1], while others are new. ✗ indicates the presence of a bug, indicates the fixing of a bug after our reporting, and ✓ indicates the fixing of a bug by previous works.
@@
On the six public-transaction clients, mpfuzz leads to the discovery of $22$ bugs, as listed in Table III, including $7$ conforming to the known DETER attacks on clients of historical versions and $15$ new bugs on the clients of the latest versions.
Other clients, including the three PBS clients and three clients on Ethereum-like networks, are mostly forks of the Geth clients, and on them, mpfuzz discovered $13$ bugs of a similar nature to those found on Geth (of the historical and latest versions), as seen in Table III.
@@
6-A Found Eviction Attacks
Exploit $XT_{1}$: Direct eviction by future transactions: In this attack, given a mempool’s initial state storing normal transactions, the attacker sends future transactions at high fees to evict the normal transactions. This exploit is essentially the DETER-X attack [1].
In practice, Geth ($\leq{}v1.11.4$), Nethermind, Besu, and EigenPhi are vulnerable under this exploit, as in Table III.
Exploit $XT_{2}$: Direct eviction by latent overdraft transactions: In this eviction attack, the attacker sends latent-overdraft transactions at high fees to evict the normal transactions initially stored in the target mempool. These transactions are sent from $k$ accounts, each of which sends $l$ transactions, denoted by $k\times{}l$. The intention is to evade the limit on the number of transactions per sender. The evasion increases the attacker cost from one pending transaction to multiple. This exploit is essentially the DETER-Z attack [1].
@@ 6-A Found Eviction Attacks
Exploit $XT_{4}$: Indirect eviction by valid-turned-overdraft transactions: This eviction attack works in two steps: 1) The attacker first sends valid transactions at high fees to evict normal transactions initially stored in the mempool. These transactions are sent from $k$ accounts, each of which sends $l$ transactions. 2) She then sends $k$ transactions; each of the transactions is of nonce $1$, at a high fee, of high value $v_{1}$, and from the same $k$ accounts in Step 1). The transactions would replace the transaction of the same sender and nonce sent in Step 1). Once the replacement is finished, they turn their child transactions into latent overdraft. Specifically, if a sender’s balance is $bal$ and the Ether value of transaction of nonce 2 is $v_{2}$, then $v_{1}$ is carefully crafted to enable turned latent overdraft, that is, $v_{1}<bal$ and $v_{1}+v_{2}>bal$.
@@
Exploit $XT_{6}$ works in three steps: 1) it first evicts the initial mempool of normal transactions with $m^{\prime}/py_{2}^{\prime}=384$ sequences, each of $16$ transactions sent from one unique sender, 2) it then sends $py_{1}^{\prime}/py_{2}^{\prime}=64$ transactions to evict the parent transactions and creates $py_{1}^{\prime}=1024$ future transactions in the mempool, 3) it sends one sequence of $5120$ transactions from one sender to evict the normal transactions sent in Step 1), and 4) it sends one transaction to evict the parent transaction in Step 3, leaving the mempool of just one valid transaction. Step 3) can succeed because the precondition (w.r.t. $py_{3}^{\prime}=5120$) of limiting transactions of the same sender does not hold.
@@
The mempool on Nethermind $<v1.18.0$ can be reversible: Minimally, suppose a two-slot mempool stores $tx_{1}$ of low fee from one sender and $tx_{2}$ of medium fee from another sender $A2$ and receives $tx_{3}$ as a child of $tx_{1}$ and of high fee. For instance, $tx_{1}\begin{bmatrix}{A}&{1}\\ {1}&{1}\end{bmatrix}$, $tx_{2}\begin{bmatrix}{B}&{1}\\ {1}&{3}\end{bmatrix}$, $tx_{3}\begin{bmatrix}{A}&{1}\\ {2}&{5}\end{bmatrix}$. The Nethermind $<v1.18.0$ mempool would admit $tx_{3}$ and evict $tx_{2}$. After that, if one resends the evicted $tx_{2}$ back to the mempool storing $tx_{1}$ and $tx_{3}$, the Nethermind $<v1.18.0$ mempool would admit $tx_{2}$ and evict $tx_{3}$, looping back to its initial state.
An attacker observing reversible mempool mounts Exploit $XT_{7}$ to evict the mempool while bringing down attack costs. On Nethermind $<v1.18.0$, the attack works in two steps: 1) She first sends $k$ transaction sequences from $k$ senders, [^4] Nethermind does not limit transactions per sender, thus $k=1$ in actual attacks. each of $l$ transactions (i.e., $k\times{}l$). In each $l$-transaction sequence, the nonce-$1$ transaction has a low fee, say $f_{1}$, while all its child transactions are of high fees, evicting normal transactions initially stored in the mempool. 2) The attacker sends $k\times(l-1)$ transactions from $k\times(l-1)$ new senders, all with fees slightly higher than $f_{1}$. The mempool would admit these transactions to the mempool by evicting all child transactions sent in Step 1), bringing down the attack costs.
6-B Found Locking Attacks
@@
Exploit $XT_{9}$: Locking mempool of no turning: Because of the risk of evicting a parent transaction that could lead to turning, certain mempool is designed to restrict eviction victims to child transactions; that is, only the transaction of maximal nonce under a given sender can be evicted. For instance, given an arriving transaction $tx$, the OpenEthereum mempool finds an existing child transaction $tx^{\prime}$ with a lower fee than $tx$ and evicts it to admit $tx$.
In Exploit $XT_{9}$, the attacker observes empty slots in the mempool (e.g., created by arriving blocks) and sends the same number of transactions to occupy them. The transactions are sent from one account where the child transaction has a higher fee $f_{1}$ than normal transactions, while all its parent transactions have minimal fees. A normal transaction that arrives subsequently is declined because $f_{1}$ is higher than the fee of a normal transaction. Because each of its parent transactions has a low fee, the mempool is locked at a low cost.
7 Attack Evaluation
7-A Evaluation of Found Eviction Exploits on Testnet
@@ 7-A1 Evaluation in Testnet
To monitor transaction propagation, we launch an independent supernode (called measurement node) in the same testnet. The measurement node runs a reconfigured Geth client where the limit of peers/neighbors is removed and the measurement node can connect to as many neighbors as possible, hence a super node. The measurement node is not (directly) connected to the attacker node. When setting up our experiments, we ran the measurement node in Goerli for seven days and found the node is stabilized at $290$ neighbors.
@@
7-B Evaluation of Found Locking Exploits on Reth
Given Exploit $XT_{8}$ found on Reth, we manually extended it to two actual exploits on unmodified Reth’s mempool, $XT_{8a}$, and $XT_{8b}$.
7-B1 Exploit Extension
@@
Actual exploit $XT_{8a}$: In $XT_{8}$, the attack cost increases with the block base fee. We manually propose a method to decrease the block base fee in Ethereum by mounting an eviction attack on the previous block. Specifically, in Ethereum (after EIP1559), given a recently produced block $i$, the base block fee of block $i+1$ is calculated dynamically as follows:
@@
Therefore, if an eviction attack can persistently lower the Gas utilization of recent blocks, the current block fees can be reduced, which decreases the costs of locking attacks on the current block. In Exploit $XT_{8a}$, we mount a series of eviction attacks first and then a locking attack against the Reth node.
Specifically, suppose in a network of a Reth node and a block validator node, the $XT_{8a}$ attacker first keeps sending eviction attacks directly to the validator node for several consecutive blocks. When observing the block base fees drop sufficiently, the attacker mounts the regular locking attack $XT_{8}$ against the Reth node.
7-B2 Attack Evaluation on Reth
@@
Figure 5: Evaluation of locking attack $XT_{8a}$ on Reth
@@
Evaluation of $XT_{8a}$: To evaluate $XT_{8a}$, we set up the workload generator that sends normal transactions to the validator Geth node. The non-validator Reth node has not yet joined the network. We let the attacker node first send eviction attacks ($XT_{6}$) directly to the victim validator for $35$ consecutive blocks until it observes the block base fees drop sufficiently to $1$ Gwei, which occurs at the $45$-th in our experiment. The Reth node then joins the network with an empty mempool. The attacker node mounts a locking attack $XT_{8}$ with transactions of price $5$ Gwei to occupy the Reth node’s mempool.